OpenStudy (anonymous):

Could someone walk me through substitution method? I have a test thursday :O x+5y=-15 5x+6y= -18

OpenStudy (perl):

in the first equation subtract 5y from both sides. this gives us x = -15 - 5y now plug this into the second equation for x 5 ( -15 - 5y ) +6y = -18 now distribute ,collect like terms and solve for y

OpenStudy (rina.r):

well their are 2 ways elimination and substitution method so u decide which one u want to apply so, Elimination works by adding or subtracting equations to get rid of one of the variables and u have to arrange the equation first for this one especially the first one which is x+5y=-15 so change this bring 15 to the left make it same like B.

OpenStudy (rina.r):

http://www.mathportal.org/algebra/solving-system-of-linear-equations/elimination-method.php this is a good example for methods try this one ok.

OpenStudy (anonymous):

Thank you! I am studying for my final on Thursday, so have to remember both methods. And of course it's comprehensive or whatever, and I did this towards the beginning of teh semester so everything is a bit foggy. @perl thank you :D I did that but somehow got lost the first time, @rena.r thank you for the link appreciate it.

OpenStudy (anonymous):

Do you want me to go through the example>

OpenStudy (anonymous):

Oh nvm

OpenStudy (anonymous):

Thanks @apple_pi :)